# Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure

### v1

@@ 3 Dynamical Model of Galaxies in Wave Dark Matter Halo
Classical dSph galaxies are known to be dominated by DM with stars being treated as the tracer particles (Gregory et al., 2019; McConnachie & Irwin, 2006; McConnachie et al., 2006; Kang & Ricotti, 2019) moving along the gravitational potential generated by the DM halo.
In this context, the corresponding velocity dispersion profile can be estimated by solving the spherically symmetric Jeans equation:
@@
where is the DM halo mass obtained by integrating the spherically symmetric density profile in Eq. (5), is the anisotropy parameter (see (Binney & Tremaine, 2008), Equation (4.61)), and is the stellar density profile as given by:
@@ 4 Simulation setup
The primary objective is to conduct a thorough comparative analysis of the stellar properties between the DM simulations and the ones stemming from both CDM and WDM simulations, spanning an extended duration down to . In the case of the WDM simulations, a particle resolution of was utilized, employing the identical hydrodynamical setup as the DM simulations. It is noteworthy that while the CDM simulation lacks any form of cutoff in its power spectrum, both the DM and WDM simulations adopt an initial power spectrum characterized by exponential suppression.
Baryons (gas and star formation) are included in the simulation, and couple to the dark matter gravitationally. The baryonic material experiences primordial and metal-line cooling, chemical enrichment, stochastic star formation with a density threshold of , supernova feedback via kinetic winds, and instantaneous uniform reionization at (Vogelsberger et al., 2013; Torrey et al., 2014; Pillepich et al., 2017; Springel et al., 2018), as has been used in the Illustris and Illustris-TNG projects. The stellar feedback model has previously been tuned to match key observables with CDM simulations (Vogelsberger et al., 2013). Whether FDM requires significant re-tuning of efficiency factors is left for future study. The subgrid model for feedback is meant to describe the effects of Type II supernovae (SNII), and uses the local star formation rate to set the mass loading of stellar winds driven by the energy available through SNII (Vogelsberger et al., 2013; Pillepich et al., 2018). The simulation also keeps track of metallicty, as described in Torrey et al. (2019). Star particles (representing a population of stars) are born with the metallicity of their surrounding interstellar gas. As star particles evolve in time, mass and metals from the aging stellar populations are returned to the interstellar medium, taking into account SN1a and SNII supernovae and AGB stars. Deposited metals in the gas evolve through passive advection with the fluid flow.
@@ 5.1 First Illustration of density profiles
Before delving into the intricacies of comparing various simulation types, we provide an initial depiction of the projected density profiles for CDM, WDM, and DM at two different redshifts. In Figure 1, we present the projected density profiles for both DM and stars. The first two rows illustrate CDM profiles, while the last two rows depict the corresponding profiles for WDM/DM. In each row, the left(right) panel presents (). Mocz et al. (2019) showed that WDM and DM exhibit remarkably similar evolutionary patterns across various redshifts, implying a resemblance in the development of large-scale structures. This stands in contrast to CDM. Consequently, in Figure 1 we exclusively focus on presenting the WDM results.
@@
Figure 1 suggests slight discrepancies between galaxies in these three distinct DM scenarios. First, in the context of CDM, filaments rapidly collapse into compact halos, whereas in the case of DM, these filaments retain a smoother configuration without giving rise to small galaxies. Second, in the CDM scenario, stars seem to exhibit greater concentration with irregular structures. Conversely, in the case of DM and WDM, a flatter shape is evident, accompanied by galaxies possessing extended stellar halos. This imparts a visual impression that stars are positioned farther away from the center. As we consider individual stellar profiles of galaxies across distinct DM scenarios, a notable finding emerges. In the DM scenario, the stellar profiles exhibit a flatter behaviour in the inner regions, in alignment with the expected pattern of stars tracing the presence of a soliton in the core. In contrast, both WDM and CDM scenarios lack such a soliton, resulting in a cuspy inner profile for these scenarios. While the primary focus of this study lies in the comparison of evolved cosmological structures between DM and CDM, it is reasonable to anticipate that the large-scale structures in WDM would bear similarities to DM. As illustrated in Mocz et al. (2019), the resulting structures are remarkably akin for WDM and DM at z=5.56, despite the more pronounced differences in their internal halo structures.
@@ 5.2 Core-halo structure
Below, we explore the presence of core-halo structures across various galaxies.
@@ 5.2.1 Signature of the core-halo structure
We start off with comparing the stellar profiles at . Figure 2 compares the stellar density profiles of the three main galaxies in the DM model at with those of the CDM model. Moving from left to right, we present the first, second, and third galaxy, respectively, as was previously pointed out in Figure 1. In the first row, a notable difference is seen between the stellar profiles in the DM scenario and CDM. This highlights how stellar profiles can vary and evolve differently in distinct DM scenarios, suggesting they could serve as a valuable tool for analyzing DM signatures in dwarf galaxies. The second row provides a closer look at the inner part of the galaxy and offers a clear test to assess the compatibility of the profiles with a core-halo structure. To this end, we plot the ideal solitonic and NFW profiles that would correspond to the inner core and the characteristic asymptotic profile, respectively. It is worth noting that only in the case of DM do the NFW and solitonic profiles intersect while still fitting the stellar profile. This is never a possibility for any CDM galaxy. This result may indicate the potential presence of a primordial core-halo stellar structure. There is also a noticeable difference in the inner profile for galaxies two and three, which aligns with the presence or absence of a soliton, resulting in a core-cuspy disagreement. Furthermore, in the case of galaxy three, where both DM and CDM exhibit a similar halo with the same slope, we can observe how the NFW profile effectively fits the entire CDM stellar profile. This is a consequence of the anticipated cusp, whereas a soliton is required for DM to establish its flat inner profile. The reason why both stellar profiles appear to coincide with a flat core for galaxy one in the inner part could be attributed to the widely suggested notion that baryonic feedback may lead to the development of a core, even in the CDM scenario (Chan et al., 2015; Pontzen & Governato, 2012). This is more plausible at this redshift for G1 than for G2 and G3, given its greater mass as the inner slope of the DM halo profiles shows a strong mass dependence (Chan et al., 2015). It is likely the same reason why the core-halo structure had sufficient time to form only in this galaxy. The vertical line in the figure represents the comoving softening length, which is 0.19 kpc in this case. Below this radius, cautious must be taken in interpreting the data.
@@
Next, we proceed with comparing the stellar profiles at between the WDM and CDM simulations, following the approach outlined in the previous section. It is worth noting that there was no available data at this redshift for DM. Despite the significant differences between WDM and DM on small scales where the core-halo structure is anticipated, their similar large-scale evolution and initial conditions could make the comparison between WDM and CDM valuable for future studies. The resulting extracted stellar profiles, shown in Figure 3, seem to reaffirm the same findings as the previous redshift for CDM. In CDM, there is no evidence or imprint of a core-halo structure identified in the halos of the three simulated galaxies. In contrast, one WDM halo exhibits a potentially compatible core-halo stellar profile, particularly in a primordial context, suggesting that they may still be in the process of forming for galaxies one and two. Furthermore, a well-defined core-halo structure is evident for the third galaxy, characterized by a notable density drop from the core to the halo, along with the distinctive change of regime point (transition point, ) marking the boundary between the inner core and the outer halo (as depicted in the lower panels of Figure 3). This is observable where the pure NFW and solitonic profiles intersect. It is interesting to observe that for WDM, the core-halo structure emerges earlier in a lower mass galaxy, such as galaxy three, in comparison to the more massive galaxy one. This structure should be established by a more pronounced core, likely a combination of a dense soliton and potentially baryonic feedback, both more prominent in heavier galaxies. Nevertheless, it is crucial to remember that WDM lacks a solitonic core. Indeed, the presence of a core-halo structure remains a topic for future analysis.
5.2.2 Stellar, dark matter and gas distribution
@@
Here we compare distinct distributions of stellar, gaseous, and dark matter profiles in all three different simulated galaxies across different DM scenarios. The aim is to extend the former comparison made only for the stellar profiles, done in Figure 2 and 3, to further include the DM and gaseous distributions as well. Figure 4 presents these profiles at . Once again, the stark contrast between the stellar and DM profiles in each case is evident, with gas profiles exhibiting a more consistent pattern across different DM models. Notably, a common feature is the presence of a core-halo transition in the DM profile from DM and gas profiles, even if it is only visibly prominent in the case of stars in G1. In comparison with DM, WDM and CDM exhibit strikingly similar cuspy stellar and DM profiles, while the gas profile resembles that of the DM case. From the plot, it is inferred that in DM model both of the stellar and gaseous profiles demonstrate similar core radii and transition points compared with the DM profile. This suggests that stars, rather than gas, may hold the key to determine the underlying DM model. Another crucial observation is the nearly identical core size and transition point in stars and gas, closely mirroring that reproduced by the DM particles. This implies a potential baryonic-dark matter connection and hints at the possibility of baryonic tracking of dark matter particles (Pozo et al., 2021; Pozo et al., 2022; Pozo et al., 2020, 2023).
@@ 5.2.3 Estimation of the core-halo parameters
We now proceed with MCMC-based DM profile fits (Eq. (7)) to the DM, as shown in Figure 5 ( Corner plot Figure 14). The core radius (), as defined in Eq. (4), plays a crucial role in characterizing the shape of the solitonic core profile. For the halo, we utilize a NFW profile with a scale radius () and the normalization factor (). The transition radius (), defining the point of density transition between the soliton and NFW profiles, is the only other free parameter. We vary this within a prior range determined by DM simulations, specifically to times (Schive et al., 2014a, b). The boson mass stands as the sole fixed parameter, maintaining a value of eV, consistent with that used for the simulations (Mocz et al., 2020). The best-fit yields a resulting profile with a total mass of , aligning well with the computed mass from Mocz et al. (2020) of for this galaxy at this redshift. In Figure 5 the blue dashed line presents the resulting profile when we conduct the analysis with the fixed total mass of , which corresponds to the total mass of that specific simulated halo (indicated by a red circle in Figure 1) at this redshift. It is worth noting that while it may not match the accuracy of the green profile with a freely varying mass, it still demonstrates excellent agreement. Furthermore, the resulting values for kpc and kpc, fall within the expected range for such dSphs, and are very close to the directly inferred soliton radius from the simulation ( kpc (Pozo et al., 2020, 2023)). This concordance between the suggested soliton size based on stellar behaviour and the actual size of the simulated DM halo soliton may hold significant implications for our understanding of how stellar behaviour is influenced by dark matter.
@@
Next, we proceed with fitting a stellar core-halo profile for another case where a core-halo shape is evident: the WDM G3 at . It is crucial to note that in principle fitting a core-halo profile to WDM, without a physical soliton, might be challenging. In this sense, we are careful not to directly link the core-halo structure to a soliton but rather to an effective profile with a transition from the core to the halo in WDM at this lower redshift. The resulting values for core radius () as well as the transition point () are summarized in Table 1, alongside other pertinent variables, with the resulting profiles depicted in Figure 6 ( Corner plot Figure 15). It is noteworthy that these parameter values fall within the expected range when compared to those extracted from observed profiles of Local Group classical dSphs. This is particularly striking considering they were analyzed at different redshifts and are still in the process of the formation. Table 1 also provides the parameter values for the two extracted galaxies and the observed dSphs in the Local Group (Pozo et al., 2023). The core radius values are in excellent agreement with the mean values obtained for such dSphs in the Local Group.
@@
The isolated galaxy (G1) from the DM simulation data (, Table 1) exhibits a core radius akin to that of two Milky Way satellite galaxies, Leo II and Sculptor. Conversely, the third galaxy from the WDM simulation data (WDM, Table 1) aligns exceptionally well with the observed cores of Sculptor and And IX. Nonetheless, there are slight disparities in the transition point, which is reasonable given the difference in redshift and the fact that these simulated galaxies are still in the process of formation. The observed increase in between and WDM could be attributed to the lower redshift of WDM, which aligns with the expectation of a larger . Notably, despite the significant redshift difference (almost 6 between the observed galaxies and the observable Local Group dwarfs), the values remain similar. This suggests that for a DM origin, the soliton is likely the first structure to form in this core-halo structure.

### v2

@@ 3 Dynamical model of galaxies in wave/fuzzy dark matter halo
Classical dSph galaxies are known to be dominated by DM with stars being treated as the tracer particles (Gregory et al. 2019; McConnachie & Irwin 2006; McConnachie et al. 2006; Kang & Ricotti 2019) moving along the gravitational potential generated by the DM halo. In this context, the corresponding velocity dispersion profile can be estimated by solving the spherically symmetric Jeans equation:
@@
Here is the DM halo mass obtained by integrating the spherically symmetric density profile in Eq. (5), describes the radial velocity dispersion, is the anisotropy parameter (see (Binney & Tremaine 2008), Equation (4.61)), and is the stellar density profile as given by
@@
We employ Eq. (7) to identify the core-halo structure in the analyzed DM stellar profiles, given its high efficiency in fitting the observed core-halo transition in local group galaxies (Pozo et al. 2024) and as it is the only profile capable of explaining this observed transition between the two regimes—inner core and outer halo—together. This approach is used exclusively for DM, as it is unsuitable for non-DM models, where the presence of a soliton is not expected (Pozo et al. 2022, 2023). For CDM–“WDM”, there is no equation that describes both regimes simultaneously, as such a structure is not expected in their DM or stellar profiles. However, for the DM profile, we can adjust eq. (5) , which defines the core-halo structure (described in Section 5.3) within the DM model, to align with what we expect for stars, as indicated by eq. (7).
For CDM/“WDM” we used the isothermal equation from Binney & Tremaine (2008) to describe the observed stellar cored profiles:
Here is the mass density at a radial distance r from the center of the galaxy. represents the central density, or the density at r = 0. is the core radius, which defines the scale at which the density profile transitions from a flat core to a declining halo.
The isothermal profile is commonly used to describe galaxies with a central region where the density remains approximately constant (the core) and a halo where the density decreases with distance, following approximately at large . However, this decrease is incompatible with the core-halo structure defined in Section 5.3 and in eq. (5), which predicts a transition point where the density should start declining according to an asymptotic profile of .
Finally, the predicted 2D stellar density profile can be projected along the line of sight, and compared with the observations, as
@@ 4 Simulation setup
The primary objective is to conduct a thorough comparative analysis of the stellar properties between the DM simulations and the ones stemming from both CDM and “WDM” simulations, spanning an extended duration down to . In the case of the “WDM” simulations, a particle resolution of was utilized, employing the identical hydrodynamical setup as the DM simulations. It is noteworthy that while the CDM simulation lacks any form of cutoff in its power spectrum, both the DM and “WDM” simulations adopt an initial power spectrum characterized by exponential suppression.
Baryons (gas and star formation) are included in the simulation, and couple to the dark matter gravitationally. The baryonic material experiences primordial and metal-line cooling, chemical enrichment, stochastic star formation with a density threshold of , supernova feedback via kinetic winds, and instantaneous uniform reionization at (Vogelsberger et al. 2013; Torrey et al. 2014; Pillepich et al. 2018a; Springel et al. 2018), as has been used in the Illustris and Illustris-TNG projects. The stellar feedback model has previously been tuned to match key observables with CDM simulations (Vogelsberger et al. 2013). Whether FDM requires significant re-tuning of efficiency factors is left for future study. The subgrid model for feedback is meant to describe the effects of Type II supernovae (SNII), and uses the local star formation rate to set the mass loading of stellar winds driven by the energy available through SNII (Vogelsberger et al. 2013; Pillepich et al. 2018b). The simulation also keeps track of metallicty, as described in Torrey et al. (2019). Star particles (representing a population of stars) are born with the metallicity of their surrounding interstellar gas. As star particles evolve in time, mass and metals from the aging stellar populations are returned to the interstellar medium, taking into account SN1a and SNII supernovae and AGB stars. Deposited metals in the gas evolve through passive advection with the fluid flow.
@@ 5.1 First illustration of density profiles
Before delving into the intricacies of comparing various simulation types, we provide an initial depiction of the projected density profiles for CDM, “WDM”, and DM at two different redshifts. In Figure 1, we present the projected density profiles for both DM and stars. The first two rows illustrate CDM profiles, while the last two rows depict the corresponding profiles for “WDM”/DM. In each row, the left(right) panel presents (). Mocz et al. (2019) showed that “WDM” and DM exhibit remarkably similar evolutionary patterns across various redshifts, implying a resemblance in the development of large-scale structures. This stands in contrast to CDM. Consequently, in Figure 1 we exclusively focus on presenting the “WDM” results.
@@
Figure 1 and 2 suggest slight discrepancies between galaxies in these three distinct DM scenarios. First, in the context of CDM, filaments rapidly collapse into compact halos, whereas in the case of DM, these filaments retain a smoother configuration without giving rise to small galaxies. Second, in the CDM scenario, stars seem to exhibit greater concentration with irregular structures. Conversely, in the case of DM and “WDM”, a flatter shape is evident, accompanied by galaxies possessing extended stellar halos. This imparts a visual impression that stars are positioned farther away from the center. As we consider individual stellar profiles of galaxies across distinct DM scenarios, a notable finding emerges. In the DM scenario, the stellar profiles exhibit a flatter behavior in the inner regions, in alignment with the expected pattern of stars tracing the presence of a soliton in the core. In contrast, both “WDM” and CDM scenarios lack such a soliton, resulting in a cuspy inner profile. While the primary focus of this study lies in the comparison of evolved cosmological structures between DM and CDM, it is reasonable to anticipate that the large-scale structures in “WDM” would bear similarities to DM. As illustrated in Mocz et al. (2019), the resulting structures are remarkably akin for “WDM” and DM at z=5.56, despite the more pronounced differences in their internal halo structures.
5.2 Stellar morphology
In this section, we compare the stellar distribution of each Galaxy across different DM simulations.
The main objective is to detect any potential imprint in the stellar structure of these Galaxies resulting from their respective dark matter models.
Since the data is limited to high values of z, it is important to note that only the results for Galaxy 1 can be considered entirely reliable, as it is the only one to have a fully virialized halo.
Figure 2 presents the stellar morphology at three projections for the isolated G1 at . The main panels in this plot display the stellar distribution for the entire Galaxy at a scale of approximately 55 kpc/h, while the additional subpanels provide a closer view of the central regions of the Galaxy (5 kpc/h). This allows for a detailed comparison between different stellar cores.
Figure 2 distinctly illustrates variations among three DM models, revealing that stars exhibit a more asymmetric structure in DM compared to a slightly symmetric shape seen in “WDM” and progressively a more symmetric shape in CDM, aligning well with theoretical predictions (Gao & Theuns 2007; Schive et al. 2020). Notably, there is a significant discernible difference between the stellar distributions in “WDM” and DM. Despite appearing similar in scale in Figure 1, stars in “WDM” are actually more flattened than in DM. Upon closer examination in the subpanels with zoomed data, stars exhibit a spherical form in the DM core, but a more elliptical shape toward the center for both CDM and “WDM”. The plot effectively highlights the disparities between the solitonic core (DM) and cuspy dark matter centers (CDM)/“WDM”. Additionally, aside from having fewer stars than CDM and “WDM”, DM has a smaller proportion of stars in the central regions (see Table 4). This suggests a more extended stellar profile around the soliton, in line with recent discoveries of halos around some studied dwarfs (Chiti et al. 2021; Collins et al. 2021). These extended halos are inherent to DM, composed of excited states above the ground state soliton, as evidenced by the NFW form, as predicted by the DM simulations (Schive et al. 2014a). This reflects the fundamentally nonrelativistic nature of DM. Moving to Figure 10, we observe “WDM”–CDM data for z 2.23. The results are comparable for both models when compared to the data from redshift z 5.56, showing an expected symmetry.
The additional subpanels in Figure 2 prominently display the predicted asymmetry for DM, a distinctive feature not observed in the symmetric distributions of stars in “WDM” and CDM models (Schive et al. 2020). This asymmetry in DM arises from the solitonic characteristic motion, which includes a random walk caused by interference on the de Broglie scale, where the spatial position of the center of the soliton changes with time, and the inherent oscillation ’jumping’ behavior (soliton oscillations) on the scale of the de Broglie wavelength or the core radius (), owing to its standing wave nature (Li et al. 2021; Chiang et al. 2021). The jumps occur over a timescale of years (Pozo et al. 2023) for a dSph with 0.3 kpc and 10 km/s, values closely matching those for this Galaxy (as referred in Tables 1 and 2). Notably, this timescale is equivalent to the stellar crossing time of the soliton, . In essence, this means that stars “sense” the movement of the soliton and respond accordingly, resulting in the observed asymmetry (Li et al. 2021; Chiang et al. 2021). We are exclusively addressing the observed comet tail in the core of the Galaxy.
@@ 5.3 Core-halo structure
Below, we explore the presence of core-halo structures across various Galaxies. Before we proceed, it is important to define some concepts for the analysis.
We define the core-halo structure as the characteristic resulting profile observed in all ideal DM dark matter models, featuring an inner flat core that can be fitted with a soliton and an outer characteristic asymptotic profile. The transition point, , is simply the radius at which this change of regime occurs. This characteristic profile has been observed in the stellar profiles of many Local Group galaxies (Pozo et al. 2024, 2023), motivating the search for such structure in the stellar profiles of the simulations presented in this work.
To analyze the compatibility of the simulated stellar profiles with the searched core-halo structure, we fit each stellar profile using a core model for the inner region (Eq. 8 for DM, and an isothermal model (Binney & Tremaine (2008), Eq. 9) for “WDM” and CDM, since Eq. 8 assumes a galaxy hosting a soliton, making it incompatible with “WDM” and CDM). For the outer halo, we use an NFW-like profile (Navarro & Frenk 1996). A clear intersection between these two profiles indicates a high level of compatibility with the investigated core-halo structure.
Since the data is limited to high values of z, it is important to note that only the results for Galaxy 1 can be considered entirely reliable, as it is the only one to have a fully virialized halo.
@@ 5.3.1 Signature of the core-halo structure
We start off with comparing the stellar profiles at . Figure 3 compares the stellar density profiles of the three main Galaxies in the DM model at with those of the CDM model. Moving from left to right, we present the first, second, and third Galaxy, respectively, as was previously pointed out in Figure 1. In this figure, a notable difference is seen between the stellar profiles in the DM scenario and CDM. This highlights how stellar profiles can vary and evolve differently in distinct DM scenarios, suggesting they could serve as a valuable tool for analyzing DM signatures in dwarf galaxies. If we take a closer look at the inner part of the Galaxy we can make a clear test to assess the compatibility of the profiles with a core-halo structure. To this end, we plot the ideal solitonic/isothermal and NFW profiles that would correspond to the inner core and the characteristic asymptotic profile, respectively. It is worth noting that only in the case of DMG1, for figure 3, do the NFW and solitonic/isothermal profiles intersect while still fitting the stellar profile. This is never a possibility for any CDM Galaxy, where the entire profile can be fitted by a single NFW or isothermal profiles. This result may indicate the potential presence of a primordial core-halo stellar structure. There is also a noticeable difference in the inner profile for Galaxies two and three, which aligns with the presence or absence of a soliton, resulting in a core-cuspy disagreement. Furthermore, in the case of Galaxy three, where both DM and CDM exhibit a similar halo with the same slope, we can observe how the NFW profile effectively fits the entire CDM stellar profile. This is a consequence of the anticipated cusp, whereas a soliton is required for DM to establish its flat inner profile. The reason why both stellar profiles appear to coincide with a flat core for Galaxy 1 in the inner part could be attributed to the widely suggested notion that baryonic feedback may lead to the development of a core, even in the cold dark matter (CDM) scenario (Chan et al. 2015; Pontzen & Governato 2012). This is more plausible at this redshift for G1 than for G2 and G3, given its greater mass, as the inner slope of the dark matter (DM) halo profiles shows a strong mass dependence (Chan et al. 2015). However, CDM G2 also requires a cored fit, albeit much smaller than the large flat inner core of DM G2. It is likely the same reason why the core-halo structure had sufficient time to form only in this Galaxy. The vertical line in the figure represents the comoving softening length, which is 0.19 kpc in this case. Below this radius, cautious must be taken in interpreting the data.
@@
Next, we proceed with comparing the stellar profiles at between the “WDM” and CDM simulations, following the approach outlined in the previous section. It is worth noting that there was no available data at this redshift for DM. Despite the significant differences between “WDM” and DM on small scales where the core-halo structure is anticipated, their similar large-scale evolution and initial conditions could make the comparison between “WDM” and CDM valuable for future studies. The resulting extracted stellar profiles, shown in Figure 4, seem to reaffirm the same findings as the previous redshift for CDM. In CDM, there is no evidence or imprint of a core-halo structure identified in the halos of the three simulated Galaxies. In contrast, one “WDM” halo exhibits a potentially compatible core-halo stellar profile, particularly in a primordial context, suggesting that they may still be in the process of forming for Galaxies one and two. Furthermore, a well-defined core-halo structure is evident for the third Galaxy, characterized by a notable density drop from the core to the halo, along with the distinctive change of regime point (transition point, ) marking the boundary between the inner core and the outer halo. This is observable where the pure NFW and cored profiles intersect. It is interesting to observe that for “WDM”, the core-halo structure emerges earlier in a lower mass galaxy, such as Galaxy three, in comparison to the more massive Galaxy one. This structure should be established by a more pronounced core, likely a combination of a dense soliton and potentially baryonic feedback, both more prominent in heavier galaxies. Nevertheless, it is crucial to remember that “WDM” lacks a solitonic core. The presence of a core-halo structure remains intriguing for “WDM” and that’s the main reason why we only proceed to evaluate it for DMG1, figure 6.
5.3.2 Stellar, dark matter, and gas distribution
@@
Here we compare distinct distributions of stellar, gaseous, and dark matter profiles in all three different simulated Galaxies across different DM scenarios. The aim is to extend the former comparison made only for the stellar profiles, done in Figure 3 and 4, to further include the DM and gaseous distributions as well. Figure 5 presents these profiles at . Once again, the stark contrast between the stellar and DM profiles in each case is evident, with gas profiles exhibiting a more consistent pattern across different DM models. Notably, a common feature is the presence of a core-halo transition in the DM profile from DM and gas profiles, even if it is only visibly prominent in the case of stars in G1. In comparison with DM, “WDM” and CDM exhibit strikingly similar cuspy DM profiles with smaller flat inner regions, while the gas profile resembles that of the DM case. From the plot, it is inferred that in DM model both of the stellar and gaseous profiles demonstrate similar core radii and transition points compared with the DM profile (for G1). This suggests that stars, rather than gas, may hold the key to determine the underlying DM model. Another crucial observation is the nearly identical core size and transition point in stars and gas, closely mirroring that reproduced by the DM particles (for DM G1). This implies a potential baryonic-dark matter connection and hints at the possibility of baryonic tracking of dark matter particles (Pozo et al. 2021, 2022, 2024, 2023). It is also important to remark that stellar profiles exhibit differences among the three DM models, with larger flat cores and more prominent declines in the DM cases compared to the smoother shapes of “WDM”–CDM. This reveals that stellar profiles evolve and behave slightly differently depending on the DM models, reinforcing their potential utility in distinguishing between the underlying DM models.
@@ 5.3.3 Estimation of the core-halo parameters
We now proceed with MCMC-based DM stellar profile fits (Eq. (7),) to the DMG1, as shown in Figure 6 ( Corner plot Figure 13). The core radius (), as defined in Eq. (4), plays a crucial role in characterizing the shape of the solitonic core profile. For the halo, we utilize a NFW profile with a scale radius () and the normalization factor (). The transition radius (), defining the point of density transition between the soliton and NFW profiles, is the only other free parameter. We vary this within a prior range determined by DM simulations, specifically to times (Schive et al. 2014a, b). The boson mass stands as the sole fixed parameter, maintaining a value of eV, consistent with that used for the simulations (Mocz et al. 2020). The best-fit yields a resulting profile with a total mass of , aligning well with the computed mass from Mocz et al. (2020) of for this Galaxy at this redshift. In Figure 6 the blue dashed line presents the resulting profile when we conduct the analysis with the fixed total mass of , which corresponds to the total mass of that specific simulated halo (indicated by a red circle in Figure 1) at this redshift. It is worth noting that while it may not match the accuracy of the green profile with a freely varying mass, it still demonstrates excellent agreement. Furthermore, the resulting values for kpc and kpc, fall within the expected range for such dSphs, and are very close to the directly inferred soliton radius from the simulation ( kpc (Pozo et al. 2024, 2023)). This concordance between the suggested soliton size based on stellar behavior and the actual size of the simulated DM halo soliton may hold significant implications for our understanding of how stellar behavior is influenced by dark matter.
@@
The resulting values for core radius () as well as the transition point () are summarized in Table 1, alongside other pertinent variables. It is noteworthy that these parameter values fall within the expected range when compared to those extracted from observed profiles of Local Group classical dSphs. This is particularly striking considering they were analyzed at different redshifts and are still in the process of the formation. Table 1 also provides the parameter values for the two extracted galaxies and the observed dSphs in the Local Group (Pozo et al. 2023). The core radius values are in excellent agreement with the mean values obtained for such dSphs in the Local Group.
@@
The isolated Galaxy (G1) from the DM simulation data (, Table 1) exhibits a core radius akin to that of two Milky Way satellite galaxies, Leo II and Sculptor. Conversely, the third Galaxy from the “WDM” simulation data (third panel of Figure 4) seems to align well with the observed cores of Sculptor and And IX. Nonetheless, there are slight disparities in the transition point, which is reasonable given the difference in redshift and the fact that these simulated Galaxies are still in the process of formation. Notably, despite the significant redshift difference (almost 6 between the observed galaxies and the observable Local Group dwarfs), the values remain similar. This suggests that for a DM origin, the soliton is likely the first structure to form in this core-halo structure.